OpenStudy (kj4uts):

A local truck rental agency advertises the following offer for a medium size truck: flat fee of $30 plus $12.95 per day plus 10 cents per mile. Determine your cost to rent such a truck for 5 days if you plan to drive it for a total of 570 miles. Express your answer in dollars accurate to the nearest cent.

OpenStudy (kj4uts):

My question that I am a little confused about are they saying $30+$12.95=$42.95 a day ($42.95*5 days = $214.75) or is it $30 once and then an addition $12.95 for each day you have it. I know that 570*.10=$57. So I think that the answer is $271.75.

OpenStudy (anonymous):

It seems like its saying $30 once + $12.95 * 5 days + $0.10 * 570 miles

OpenStudy (kj4uts):

$30 $12.95*5=$64.75 $0.10*57=$57 =$151.75 Ok I see I guess the $30 is a one time fee (flat) one question do you still include the $12.95 the first day?

OpenStudy (anonymous):

Yes I read it as including $12.95 for the first day. I'd go with $151.75

OpenStudy (kj4uts):

Ok thanks again :)

OpenStudy (anonymous):

you're welcome